Grade 10Electromagnetic Induction

An alpha particle is projected vertically upward with a speed of 3.0 × 10^4 km/s in a region where a magnetic field of magnitude 1.0 T exists in the direction south to north. Find the magnetic force that acts on the α-particle.

Hrishant Goswami
12 Years agoGrade 10

1 Answer

Jitender Pal
12 Years ago
Sol. q = 2 ×1.6 × 10^–19 C, υ = 3 × 10^4 km/s = 3 × 10^7 m/s B = 1 T, F = qBυ = 2 × 1.6 × 10^–19 × 3 × 10^7 × 1 = 9.6 10^–12 N. towards west.